# Causal Interventions Reveal Shared Structure Across English Filler–Gap Constructions

### v1

@@ Results
We further find evidence supporting our hypothesis that linguistic similarity aids generalization between constructions. Our regression reveals significant, positive effects for filler type at the filler and the positions, inversion of the head daughter and nature of the fronted element at the filler, the, and np positions, and syntactic category of the mother at all positions.
Figure 5: max odds ±$\pm 1$ standard error, by position for interventions (1) trained and evaluated on multi-clause variants, (2) trained on single-clause variants and evaluated on multi-clause variants, and (3-4) controls.
@@ Discussion
These results paint a clear picture of filler–gap generalization in LMs. Frequent constructions are encountered at a high-enough rate during training so as to drive the development of robust mechanisms to process them. Less frequent constructions are not encountered enough for stand-alone, robust, processing mechanisms to form. Instead, their processing relies on the mechanisms of more frequent, linguistically similar constructions.
These analyses shed light on the anomalous results. For instance, we observed a low in-degree AUC for the low-frequency construction topicalization. Topicalization is linguistically dissimilar to higher-frequency constructions, being the only construction with a phrasal element at its filler site, and it generally shares very few linguistic features with more frequent constructions. In this light, its low in-degree AUC is not surprising, especially when compared to pseudoclefts, which much more closely resemble higher-frequency constructions (especially wh-questions).
@@
Set-Up
We evaluate the interventions trained at each position of the single-source variants on the corresponding positions in the matrix and embedded clause of the same construction’s multi-clause template. We compare the by-position max odds values to the corresponding values of interventions trained and evaluated on the multi-clause variants.
@@ Results
Our results are displayed in Figure 5. Our single-clause mechanisms show above-chance max odds at the filler through np1 positions of the matrix clause, before dropping off at the verb1 through the2 positions, and then slowly rebounding as we move towards the final verb2.
These results make sense when we consider the relative sentential structures of single-clause and multi-clause sentences, and the auto-regressive nature of the LMs we study. Primarily, the first three positions of a multi-clause sentence – that is, filler, the1, and np1 – are indistinguishable from the first three positions of a singular-clause sentence. As such, we would expect an auto-regressive LM, processing from left-to-right, to not be aware that it is processing an embedded clause until it reached the verb1 position. Until then, it will use the same mechanisms it would to process a sentence with a single clause. This is reflected in the strong generalization through these first three positions.
@@ 7 Conclusion
Long-held views in linguistics suggest that there should be common processing characteristics across diverse filler–gap constructions. We found this to be the case: we were able to transfer the filler–gap property across neural representations of different filler–gap constructions, suggesting that neural models rely on similar representations across distinct constructions.
This transfer is not entirely uncomplicated, however: transfer was stronger when animacy matched and less strong when animacy did not match even within a construction. This was true even though animacy is not a key part of the usual account of English filler–gap constructions. We also found that some constructions were stronger sources of filler–gap transfer than others — and others were stronger sinks. Finally, we found that transfer across main and embedded clauses was not strong. Taken together, these results show how mechanistic analysis of LLMs can provide novel insights into the nature of syntactic structures.
@@ 8 Limitations
Our work is primarily an attempt to show that LLMs can be useful tools for pushing linguistic theory forward. This brings with it specific theoretical presuppositions that are worth articulating to avoid a suggestion that there is scientific consensus where there is not.
Our investigation is oriented toward finding evidence of modular structure in LLMs. However, it is not a settled question what constitutes rule-like or systematic linguistic behavior in neural systems (Geiger et al., 2024; Nefdt, 2023; Buckner, 2024; Futrell and Mahowald, 2025). How causally systematic should a syntactic behavior be for it to be rule-like? One reading of our results would be that our causal interventions capture human filler–gap behavior but noisily (e.g., imperfect transfer across constructions, less transfer when animacy differs).
This is possible, but another reasonable interpretation is that the relevant constructs are also fuzzy in humans. Despite a historical proclivity for rules, nearly all syntactic theories allow for numerous exceptions, and human behavior itself is variable and subject to errors. As such, the questions we ask regarding the rule-like nature of LLMs extend beyond such models, becoming broader questions about human processing and behavior. Our findings alone cannot adjudicate these questions, though.
We also note that our results are only in English. It would be valuable to extend to other languages, particularly those with typologically different filler-gap patterns.
We relied here on templatically generated sentences, which are known to differ in systematic ways from naturally occuring sentences. We would like to extend this work to naturalistic sentences, but doing so is challenging because of the strong constraint that we have matched pairs.
@@ Acknowledgments
We would like to thank Qing Yao and, more broadly, the computational linguistics research group at UT Austin for their helpful conversations regarding this project. We further acknowledge funding from NSF CAREER grant 2339729 (to Kyle Mahowald).
@@ Appendix A Construction Templates
We provide templates and examples for our single-clause inanimate extraction (Table 3), multi-clause animate extraction (Table 4), and multi-clause inanimate extractions (Table 5). In these tables, we use the shorthand demonstrated in Table 2 to refer to our constructions.
@@
Table 2: Abbreviations for syntactic constructions in Tables 3, 4 and 5.
@@ Appendix B Training and Evaluation Details
We access the pythia models used in this study through the transformers python package (Wolf et al., 2020). For each construction, we build out training sets as described by Arora et al. (2024), sampling 200 sentences to form the basis of our training set, before adding each sentence’s minimal pair, resulting in perfectly balanced training sets. To train DAS, we use the pyvene library (Wu et al., 2024) and follow the hyperparameters used by Arora et al. (2024).
Our evaluation sets for the pythia-1.4b models used in the main text consist of 400 sentences, with odds at each position-layer pair averaged across all evaluation sentences. For the other model variants evaluated (pythia-2.8b and pythia-6.9b) we use evaluation sets of 96 sentences due to computational constraints, noting that this is still larger than the prescribed size of 50 evaluation sentences from Arora et al. (2024). We also ensure that the intersection of train sets and evaluation sets is empty, so as to not bias our evaluations. Our training and evaluation ran on 2 NVIDIA A40 GPUs. For one model size, training totaled $\sim$12 hours, and evaluation $\sim$250 hours.
[…]
(1 + in__\__train__\__set * same__\__animacy | to) +
in__\__train__\__set * same__\__animacy
@@
[…]
match__\__embedded__\__under + match__\__discourse__\__fronted || from ) +
(1 + match__\__filler__\__class + match__\__inversion +
match__\__embedded__\__under + match__\__discourse__\__fronted || to ) +
match__\__filler__\__class + match__\__inversion +
match__\__embedded__\__under + match__\__discourse__\__fronted
@@
Construction Filler Class Inverted Clause Embedding Item Fronting for Discourse
Embedded Wh-Q Wh-Item False Verb Phrase False
Matrix Wh-Q Wh-Item True N/A False
RRC Wh-Item False Noun Phrase False
Cleft Null-Element False Verb Phrase True
Pseudocleft Wh-Item False N/A True
Topicalization Phrasal Element False N/A True
Table 6: Parameter values for each filler–gap construction.
@@ C.1 Experiment 1 Regression
In the leave-one-out setting, we fit a linear mixed effects model at each position with our dependent variable as the max odds at each training-set, and evaluation-set pair. We treat the training-set and evaluation-set as random effects, and indicator variables for whether the evaluation-set comprises a construction in the training-set and whether the evaluation-set has the same animacy as the training-set as fixed effects. We also include a term to investigate their interaction. As per Barr et al. (2013), we include maximal random effect slope structures. Our full regression model is reported in Figure 6, which we fit to obtain the reported $\beta$ coefficients, and corresponding p-values.
Indicator variables are codified such that if the evaluated construction is in the training-set, in__\__train__\__set = 1 with in__\__train__\__set = -1 otherwise. Similarly, if the evaluated construction’s animacy matches that of the training conditions, same__\__animacy = 1 with same__\__animacy = -1 otherwise. The full results of this regression can be found in Table 7. Note: In this setting, the construction__\__from variable denotes the held-out construction.
Term $\beta$${}_{\textsc{filler}}$ $\beta$${}_{\textsc{the}}$ $\beta$${}_{\textsc{np}}$ $\beta$${}_{\textsc{verb}}$
@@
[…]
@@ C.2 Experiment 2 Regression
In the single-construction setting, we fit a linear mixed effects model at each position with our dependent variable as the max odds at each training-set and evaluation-set pair. We treat the training-set and evaluation-set as random effects. Our mixed effects comprise indicator variable denoting whether the training construction and the evaluation construction match in our proposed filler–gap parameters of variation. A full breakdown of these parameters of variation and how they apply to our constructions of interest can be seen in Table 6. The resulting indicator variables take a value of 1 if the construction in the trainset and the construction in the evaluation set match for that given parameter, and -1 otherwise. We include maximal random effect slope structures, excluding correlations to help convergence, as per Barr et al. (2013).
Our resulting regression model is reported in in Figure 7, which we fit to obtain the reported $\beta$ coefficients, and corresponding p-values (Table 8).
Term $\beta$${}_{\textsc{filler}}$ $\beta$${}_{\textsc{the}}$ $\beta$${}_{\textsc{np}}$ $\beta$${}_{\textsc{verb}}$
(Intercept) 1.15*** 1.96*** 1.32*** 7.12***
match_filler_class 0.75*** 1.06** 0.28 0.53
match_inversion 0.38** 0.51** 0.40** 0.06
match_embedded_under 0.85*** 1.05** 0.54** 2.06**
match_discourse_fronted 0.30** 0.36* 0.34*** 0.32
[…]
@@ Appendix D Frequencies
To calculate frequencies, we use the English-EWT Universal Dependencies dataset (De Marneffe et al., 2021; Nivre et al., 2020; Silveira et al., 2014). It is sourced from the English Web Treebank, a corpus which totals 16,622 sentences scraped from the web. We parse the train, test, and dev connlu associated files searching for dependency relations denoting each of our given constructions. We do not differentiate between our two classes of embedded wh-questions, as the lexically defined constraint would have likely yielded a non-exhaustive extraction of all possible sentences. Instead we calculate a generic total for embedded wh-questions, and share this count among both of them. We present the final counts in Table 9.
@@ Appendix E Experiment 1: Supplementary Information
A by-position aggregation figure for the multi-clause variant is in Figure 8, complementing Figure 2. An extended version of the mechanistic plots in Figure 3, including controls, appears in Figure 9, with a multi-clause counterpart shown in Figure 10.
@@ Appendix F Experiment 2: Supplementary Information
We report raw bar charts for AUCs of in-degree and out-degree centrality across single- and multi-clause settings (Figures 11, 12, 13 and 14).
@@ Appendix G Experiment 3: Supplementary Information
We also provide mechanistic heatmaps for our cross-clausal generalization experiments. They can be found in Figure 15.
@@
Appendix H Duplication with Other Model Sizes
We duplicate these experiments with other model sizes, namely pythia-2.8b and pythia-6.9b. Below, we report these results.
@@ H.1 Experiment 1
We provide the aggregation figures across positions – single (Figure 16) and multi-clause (Figure 17) variants. We note that we find significant differences in the same positions as with the pythia-1.4b models. We provide regression results in Table 10.
@@
Figure 16: Top: pythia-2.8b and bottom: pythia-6.9b average normalized max odds across positions in the single-clause variants, ±$\pm 1$ standard error. Normalization fixes the “Same Animacy, In Train Set” condition at 1.00.
@@
Figure 17: Top: pythia-2.8b and bottom: pythia-6.9b average normalized max odds across positions in the single-clause variants, ±$\pm 1$ standard error. Normalization fixes the “Same Animacy, In Train Set” condition at 1.00.
Term $\beta$${}_{\textsc{filler}}$ $\beta$${}_{\textsc{the}}$ $\beta$${}_{\textsc{np}}$ $\beta$${}_{\textsc{verb}}$
@@
H.2 Experiment 2
For experiment 2, we provide scatter plots in Figure 18 and regression results in Table 11.
Term $\beta$${}_{\textsc{filler}}$ $\beta$${}_{\textsc{the}}$ $\beta$${}_{\textsc{np}}$ $\beta$${}_{\textsc{verb}}$
@@ H.2 Experiment 2
[…]

### v2

@@ Results
[…]
(Intercept) 1.15*** 1.96*** 1.32*** 7.12***
filler 0.75*** 1.06** 0.28 0.53
inverted 0.38** 0.51** 0.40** 0.06
embed 0.85*** 1.05** 0.54** 2.06**
front 0.30** 0.36* 0.34*** 0.32
Table 2: Experiment 2 Regression Results. * denotes $p<.05$, ** denotes $p<.01$, and *** denotes $p<.001$. The dependent variable is the max odds The coefficients correspond to the linguistic variables of interest shown in Table 1. Generalization tends to be significantly greater when linguistic properties are shared.
We further find evidence supporting our hypothesis that linguistic similarity aids generalization between constructions. Our regression (Table 2) reveals significant, positive effects for filler type at the filler and the positions, inversion of the head child and nature of the fronted element at the filler, the, and np positions, and syntactic category of the parent at all positions. Furthermore, Figure 5 demonstrates convergent results from PCA, with linguistically related constructions generally clustering along the principal components. For instance, animate and inanimate forms of the same construction tend to cluster together, and cleft and topicalization constructions tend to cluster together.
@@ Discussion
These results paint a clear picture of filler–gap generalization in LMs. Frequent constructions are encountered at a high-enough rate during training to drive the development of robust mechanisms to process them. Less frequent constructions are not encountered enough for stand-alone, robust processing mechanisms to form. Instead, their processing relies on the mechanisms of more frequent, linguistically similar constructions.
Further linguistic analyses reveal effects beyond frequency. For instance, we observed a low in-degree AUC for the low-frequency construction topicalization. Topicalization is linguistically dissimilar to higher-frequency constructions, being the only construction with a phrasal element at its filler site, and it generally shares very few linguistic features with more frequent constructions. In this light, its low in-degree AUC is not surprising, especially when compared to pseudo-clefts, which much more closely resemble higher-frequency constructions (especially wh-questions).
@@
Figure 6: max odds $\pm 1$ standard error, by position for interventions (1) trained and evaluated on multi-clause variants, (2) trained on single-clause variants and evaluated on multi-clause variants, and (3–4) controls.
@@
Setup
We evaluate the interventions trained at each position of the single-clause variants on the corresponding positions in the matrix and embedded clause of the same construction’s multi-clause template. We compare the by-position max odds values to the corresponding values of interventions trained and evaluated on the multi-clause variants.
@@ Results
Our results are displayed in Figure˜6. Overall, while we see robust transfer between embedded clauses, we see little meaningful transfer from matrix to embedded clauses.
Our single-clause mechanisms show above-chance max odds at the filler through np1 positions of the matrix clause, before dropping off at the verb1 through the2 positions, and then slowly rebounding as we move towards the final verb2.
These results make sense when we consider the relative sentential structures of single-clause and multi-clause sentences, and the auto-regressive nature of the LMs we study. The first three positions of a multi-clause sentence – that is, filler, the1, and np1 – are indistinguishable from the first three positions of a singular-clause sentence. As such, we would expect an auto-regressive LM, processing from left-to-right, to not be aware that it is processing an embedded clause until it reached the verb1 position. Until then, it will use the same mechanisms it would to process a sentence with a single clause. This is reflected in the strong generalization through these first three positions.
@@
Discussion
While many syntactic theories posit that filler-gap structures are processed uniformly across contexts, our findings suggest that, in LMs, filler-gap constructions are handled by different mechanisms in matrix and embedded clauses.
@@ 7 Conclusion
Long-held views in linguistics suggest that there should be common processing characteristics across diverse English filler–gap constructions. We found this largely to be the case for LMs: we were able to transfer the filler–gap property across neural representations of different filler–gap constructions, suggesting that neural models rely on similar representations across distinct constructions.
Moreover, our analyses suggest that the strength of the transfer is mediated by linguistically interesting properties. We see a significant, positive boost in generalization for constructions matching in filler type, for constructions with similar verb inversion patterns, for constructions that match in whether they involve information-structural fronting, and for constructions in which the relevant syntactic parents share a syntactic category.
The transfer effects are not uncomplicated, though. The observed structural effects are accompanied by frequency and animacy effects. The processing mechanisms of more frequent constructions support the processing of less frequent constructions. And, even across constructions which are syntactically identical but differ in animacy, transfer is weaker than when animacy matches. This was true even though animacy and frequency are not a key part of the usual syntactic account of filler–gap constructions. We also found transfer between embedded and matrix clauses to be weak.
These findings point to linguistically interesting hypotheses about the factors governing constructional similarity – hypotheses that could directly inform future linguistic research by, for instance, exploring whether humans show greater priming effects across filler–gap constructions that share these relevant properties. We argue that mechanistic analysis of LMs can provide novel insights into the nature of syntactic structures.
@@ 8 Limitations
Our work is primarily an attempt to show that LMs can be useful tools for pushing linguistic theory forward. This brings with it specific theoretical presuppositions that are worth articulating to avoid a suggestion that there is scientific consensus where there is not.
Our investigation is oriented toward finding evidence of modular structure in LMs. However, it is not a settled question what constitutes rule-like or systematic linguistic behavior in neural systems (Nefdt, 2023; Geiger et al., 2024; Buckner, 2024; Futrell and Mahowald, 2025). How causally systematic should a syntactic behavior be for it to be rule-like? One reading of our results would be that our causal interventions capture human filler–gap behavior but noisily (e.g., imperfect transfer across constructions, less transfer when animacy differs).
This is possible, but another reasonable interpretation is that the relevant constructs are also fuzzy in humans. Despite a historical proclivity for rules, nearly all syntactic theories allow for numerous exceptions, and human behavior itself is variable and subject to errors. As such, the questions we ask regarding the rule-like nature of LMs extend beyond such models, becoming broader questions about human processing and behavior. Our findings alone cannot adjudicate these questions, though.
Further, while we provide evidence in Section˜5 that the frequency of a given construction plays a large role in its strength as a source, we do not preclude that this is the only factor driving source strength. Specifically, we do not rule out that the inductive biases present in Transformer-based LMs may inherently process certain constructions better than others. However, our study is not designed in a manner such that we can address this, and we leave this as a direction for future work.
We also note that our results are only in English. It would be valuable to extend them to other languages, particularly those with typologically different filler–gap patterns.
We relied here on templatically generated sentences, which are known to differ in systematic ways from naturally occurring sentences. We would like to extend this work to naturalistic sentences, but doing so is challenging because of the strong constraint that we have matched pairs.
@@ Acknowledgments
We would like to thank Qing Yao and, more broadly, the whole computational linguistics research group at UT Austin for their helpful conversations regarding this project. We thank audiences at Saarland University, Edinburgh University, and Dagstuhl Seminar 25301 for helpful comments. We acknowledge funding from NSF CAREER grant 2339729 to Kyle Mahowald and from Google and Open Philanthropy to Christopher Potts.
@@ Appendix A Construction Templates
We provide templates and examples for our single-clause inanimate extraction (Table˜4), multi-clause animate extraction (Table˜5), and multi-clause inanimate extractions (Table˜6). In these tables, we use the shorthand demonstrated in Table˜3 to refer to our constructions.
@@
Table 3: Abbreviations for syntactic constructions in Tables˜4, 5 and 6.
@@ Appendix B Training and Evaluation Details
We access the pythia models used in this study through the transformers python package (Wolf et al., 2020). To train DAS, we use the pyvene library (Wu et al., 2024) and follow the hyperparameters used by Arora et al. (2024).
Our evaluation sets for the pythia-1.4b models consist of 400 sentences, with odds at each position-layer pair averaged across all evaluation sentences. For the other model variants evaluated (pythia-2.8b and pythia-6.9b) we use evaluation sets of 96 sentences due to computational constraints, noting that this is still larger than the prescribed evaluation size of 50 sentences from Arora et al. (2024). We ensure that the intersect of train sets and evaluation sets is empty, so as to not bias our evaluations. Our training and evaluation ran on 2 NVIDIA A40 GPUs. For one model size, training totaled $\approx$12 hours, and evaluation $\approx$250 hours.
$\texttt{model <- lmer(max$\_$odds)}\sim$ (1 + in_\_train_\_set * same_\_animacy | from) +
(1 + in_\_train_\_set * same_\_animacy | to) +
in_\_train_\_set * same_\_animacy
@@
$\texttt{model <- lmer(max$\_$odds)}\sim$ (1 + match_\_filler_\_class + match_\_inversion +
match_\_embedded_\_under + match_\_discourse_\_fronted || from ) +
(1 + match_\_filler_\_class + match_\_inversion +
match_\_embedded_\_under + match_\_discourse_\_fronted || to ) +
match_\_filler_\_class + match_\_inversion +
match_\_embedded_\_under + match_\_discourse_\_fronted
@@ C.1 Experiment 1 Regression
In the leave-one-out setting, we fit a linear mixed-effects model at each position with our dependent variable as the max odds at each training – evaluation-set pair. We treat the training-set and evaluation-set as random effects, with fixed effects comprising indicator variables for whether the constructions in the training-set and evaluation-set match and whether animacy of the training-set and evaluation-set match. We also include a term for their interaction. As per Barr et al. (2013), we include maximal random effect slope structures. Our full regression model is as in Figure˜7, which we fit to obtain the reported $\beta$ coefficients, and corresponding p-values.
Indicator variables are codified such that if the evaluated construction is in the training-set, in_\_train_\_set = 1 with in_\_train_\_set = -1 otherwise. Similarly, if the evaluated construction’s animacy matches that of the training conditions, same_\_animacy = 1 with same_\_animacy = -1 otherwise. Table 7 shows full regression results. Note: In this setting, the construction_\_from variable denotes the held-out construction.
[…]
@@
Table 7: Experiment 1 Regression Results. * denotes $p<.05$, ** denotes $p<.01$, and *** denotes $p<.001$. The dependent variable is the Max Odds. Coefficients correspond to relationship between the training set and evaluation set for a given intervention. Generalization tends to be stronger when the evaluated constructions are in the training set and match in animacy.
@@ C.2 Experiment 2 Regression
In the single-construction setting, we fit a linear mixed-effects model at each position with our dependent variable as the max odds at each training-set and evaluation-set pair. We treat the training-set and evaluation-set as random effects. Our mixed-effects comprise indicator variables denoting whether the training construction and the evaluation construction match in our proposed filler–gap parameters of variation. A full breakdown of these parameters of variation and how they apply to our constructions of interest can be seen in Table 1. The resulting indicator variables take a value of 1 if the construction in the trainset and the construction in the evaluation set match for that given parameter, and -1 otherwise. We include maximal random effect slope structures, excluding correlations to help convergence, as per Barr et al. (2013).
Our resulting regression model is reported in in Figure˜8, which we fit to obtain the reported $\beta$ coefficients, and corresponding p-values (Table 2).
@@ Appendix D Frequencies
To calculate frequencies, we use the English-EWT Universal Dependencies dataset (De Marneffe et al., 2021; Nivre et al., 2020; Silveira et al., 2014). It is sourced from the English Web Treebank, a corpus which totals 16,622 sentences scraped from the web. We parse the train, test, and dev CoNLL-U associated files searching for dependency relations denoting each of our given constructions. We do not differentiate between our two classes of embedded wh-questions, as the lexically defined constraint would have likely yielded a non-exhaustive extraction of all possible sentences. Instead we calculate a generic total for embedded wh-questions, and share this count among both of them. We present the final counts in Table 8.
@@ Appendix E Experiment 1: Supplementary Information
A by-position aggregation figure for the multi-clause variant is in Figure˜9, complementing Figure˜2. An extended version of the mechanistic plots in Figure˜3, including controls, appears in Figure˜10, with a multi-clause counterpart shown in Figure˜11.
@@ Appendix F Experiment 2: Supplementary Information
We report raw bar charts for AUCs of in-degree and out-degree centrality across single- and multi-clause settings (Figures˜12, 13, 14 and 15).
@@ Appendix G Experiment 3: Supplementary Information
We also provide mechanistic heatmaps for our cross-clausal generalization experiments. They can be found in Figure˜16.
@@
Appendix H Replication with Other Model Sizes
We replicate these experiments with other model sizes, namely pythia-2.8b and pythia-6.9b. Below, we report these results.
@@ H.1 Experiment 1
We provide the aggregation figures across positions – single (Figure˜17) and multi-clause (Figure˜18) variants. We note that we find significant differences in the same positions as with the pythia-1.4b models. We provide regression results in Table˜9.
@@
Figure 17: Top: pythia-2.8b and bottom: pythia-6.9b average normalized max odds across positions in the single-clause variants, $\pm 1$ standard error. Normalization fixes the “Same Animacy, In Train Set” condition at 1.00.
@@
Figure 18: Top: pythia-2.8b and bottom: pythia-6.9b average normalized max odds across positions in the single-clause variants, $\pm 1$ standard error. Normalization fixes the “Same Animacy, In Train Set” condition at 1.00.
[…]
@@
Table 9: Experiment 1 Regression Results for pythia-2.8b and pythia-6.9b. * denotes $p<.05$, ** denotes $p<.01$, and *** denotes $p<.001$.
H.2 Experiment 2
For experiment 2, we provide scatter plots in Figure˜19, regression results in Table 10, and PCA plots in Figure˜20.
@@ H.2 Experiment 2
Table 10: Experiment 2 Regression Results for pythia-2.8b and pythia-6.9b. * denotes $p<.05$, ** denotes $p<.01$, and *** denotes $p<.001$ .